### v1

# A Micro-Macro Decomposition-Based Asymptotic-Preserving Random Feature Method for Multiscale Radiative Transfer Equations

Abstract
This paper introduces the Asymptotic-Preserving Random Feature Method (APRFM) for the efficient resolution of multiscale radiative transfer equations. The APRFM effectively addresses the challenges posed by stiffness and multiscale characteristics inherent in radiative transfer equations through the application of a micro-macro decomposition strategy. This approach decomposes the distribution function into equilibrium and non-equilibrium components, allowing for the approximation of both parts through the random feature method (RFM) within a least squares minimization framework. The proposed method exhibits remarkable robustness across different scales and achieves high accuracy with fewer degrees of freedom and collocation points than the vanilla RFM. Additionally, compared to the deep neural network-based method, our approach offers significant advantages in terms of parameter efficiency and computational speed. These benefits have been substantiated through numerous numerical experiments conducted on both one- and two-dimensional problems.
1 Introduction
The radiative transfer equation (RTE) is the governing equation that models the propagation and interactions of radiation or particles within participating media [1, 2]. It is a fundamental integro-differential equation in various fields, including astrophysics, radiative transfer [3], neutron transport [4] and optical tomography [5, 6, 7, 8], etc. In recent years, there has been significant interest in devising accurate and efficient methods for solving the multiscale radiative transfer equation. The primary bottleneck in numerically resolving the radiative transfer equation stems from the high dimensionality of phase space, stiffness of collision terms, and multiscale features, among others. Various numerical methods have been developed in the field of computational methods for the radiative transfer equation, which can generally be categorized into two classes: deterministic methods and stochastic simulation methods. One of the most popular deterministic methods is the discrete ordinates/velocity method (DOM/DVM) [4, 9], sometimes referred to as the $S_{N}$ method. The DOM discretizes the angular variable and solves the RTE along the discrete directions. Spherical harmonics methods possess the advantage of rotational invariance and are widely used in solving the radiative transfer equation [10, 11]. Note that the scale parameter (Knudsen number) in the radiative transfer equation can vary significantly, ranging from the kinetic regime to the diffusive regime, therefore numerical methods should be able to handle the multiscale nature of the radiative transfer equation. To tackle this challenge, two major categories of methods have been developed, namely, the domain decomposition-based methods [12] and the asymptotic-preserving schemes [13]. Domain decomposition-based methods decompose the domain into different regions, where different differential equations are solved in each region with an interface condition to couple them. The asymptotic-preserving schemes aim to design numerical methods that are uniformly stable and accurate, regardless of the scale parameter. For stochastic simulation methods, the direct simulation Monte Carlo (DSMC) method is widely used for solving the radiative transfer equation [14]. There are also some references on solving the radiative transfer equation, interested readers can refer to [15, 16, 17, 18, 19, 20, 21] for more details.
In the past few years, deep learning methods have shown great potential in solving high-dimensional and complex PDE problems, due to their strong fitting ability and generalization ability. Researchers have developed some deep learning-based methods for solving the radiative transfer equation and kinetic equations [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. Although deep learning methods have achieved great success in solving PDEs, they still face challenges in terms of interpretability, generalization, and computational efficiency. For example, the training time for deep learning methods can be very long, and the accuracy of the solution may be less than satisfactory. The random feature method (RFM) bridges the gap between deep learning methods and traditional numerical methods, and is effective in solving various PDEs [39, 40, 41, 42]. The RFM approximates the solution of the PDE by a linear combination of random feature functions and Partition of Unity (PoU) functions, and the coefficients are determined by a least squares minimization problem. In particular, the random feature functions are constructed by two-layer neural networks with fixed parameters in the hidden layers and the PoU functions are constructed by tensor product of univariate functions. Benefiting from the construction of the random feature functions and least squares minimization problem, the RFM can achieve high accuracy and efficiency in solving PDEs.
In this paper, we start by introducing the RFM for solving the radiative transfer equation and find that the vanilla random feature method has tremendous difficulty in resolving small scales. To address this issue, we propose the Asymptotic-Preserving Random Feature Method (APRFM) for solving the multiscale radiative transfer equation. The APRFM is designed to effectively handle the multiscale nature of the radiative transfer equation by utilizing a micro-macro decomposition approach. Our method can approximate the solution of the radiative transfer equation by decomposing the distribution function into equilibrium and non-equilibrium components and approximating both parts through the RFM within a least squares minimization framework. The proposed method demonstrates superior robustness across varying scales compared to the vanilla RFM and is more efficient than the previous deep learning methods.
The rest of the paper is organized as follows. In Section 2, we gave a brief introduction to the radiative transfer equation and the random feature method. Besides, we demonstrate the difficulty of the vanilla random feature method in resolving small scales. Section 3 is the main part of the paper, where we propose the Asymptotic-Preserving Random Feature Method. In Section 4, we present numerical results for both one- and two-dimensional problems to validate the effectiveness of our method. Finally, we conclude the paper in the last section.
2 Preliminaries
2.1 The radiative transfer equation
We consider the scaled form of the stationary radiative transfer equation on a bounded Lipschitz domain $D\subset\mathbb{R}^{d}$ as follows:
$\bm{v}\cdot\nabla_{\bm{x}}f(\bm{x},\bm{v})=\frac{\sigma_{s}(\bm{x})}{\varepsilon}\mathcal{L}f(\bm{x},\bm{v})-\varepsilon\sigma_{a}(\bm{x})f(\bm{x},\bm{v})+\varepsilon Q(\bm{x},\bm{v}),\;(\bm{x},\bm{v})\in D\times\mathbb{S}^{d-1},$ (1)
where $f(\bm{x},\bm{v})$ denotes the distribution function of particles at space position $\bm{x}\in D$ and velocity direction $\bm{v}\in\mathbb{S}^{d-1}$, and $Q(\bm{x},\bm{v})$ represents the source function. The non-negative functions $\sigma_{s}(\bm{x})$ and $\sigma_{a}(\bm{x})$ correspond to the scattering coefficient and absorption coefficient, respectively. The dimensionless parameter $\varepsilon>0$, referred to as the Knudsen number, captures the ratio of the mean free path to the characteristic length of the domain. In the context of the multiscale problem, the Knudsen number $\varepsilon$ spans magnitudes from $\mathcal{O}(1)$ (kinetic regime) to $\varepsilon\ll 1$ (diffusive regime). A smaller Knudsen number $\varepsilon$ implies more frequent collisions between particles. The operator $\mathcal{L}$ is defined as:
$\mathcal{L}f=\frac{1}{|\mathbb{S}^{d-1}|}\int_{\mathbb{S}^{d-1}}k(\bm{v},\bm{v}^{\prime})(f(\bm{v}^{\prime})-f(\bm{v}))\mathop{}\!\mathrm{d}\bm{v}^{\prime},$ (2)
where $k:\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}\to\mathbb{R}$ is a non-negative kernel. Here, for convenience, we employ the notation
$\left\langle h\right\rangle:=\frac{1}{|\mathbb{S}^{d-1}|}\int_{\mathbb{S}^{d-1}}h(\bm{v}^{\prime})\mathop{}\!\mathrm{d}\bm{v}^{\prime},$ (3)
which denotes the velocity angular average of the function $h$ over the unit sphere $\mathbb{S}^{d-1}$. Next, we make some important assumptions [43] on the operator $\mathcal{L}$ for the well-posedness of the radiative transfer equation (1):
  • $\left\langle\mathcal{L}f\right\rangle=0,\;\forall f(\bm{v})\in L^{2}(\mathbb{S}^{d-1})$.
  • $\mathcal{L}$ is a self-adjoint and non-positive operator in $L^{2}(\mathbb{S}^{d-1})$.
  • The null space of $\mathcal{L}$ is $\{f\in L^{2}(D\times\mathbb{S}^{d-1}):f=\left\langle f\right\rangle\}$.
In this paper, we focus on the inflow boundary:
$\Gamma_{-}:=\{(\bm{x},\bm{v})\in\partial D\times\mathbb{S}^{d-1}|\ \bm{v}\cdot\bm{n}(\bm{x})<0\},$ (4)
where $\bm{n}(\bm{x})$ represents the outward normal vector to the boundary $\partial D$.
In the diffusive limit, the radiative transfer equation (1) can be approximated by the elliptic equation [44, 27]:
$\left\langle\bm{v}\cdot\nabla_{\bm{x}}\mathcal{L}^{-1}\left(\frac{1}{\sigma_{s}(\bm{x})}\bm{v}\cdot\nabla_{\bm{x}}\rho(\bm{x})\right)\right\rangle=-\sigma_{a}(\bm{x})\rho(\bm{x})+\left\langle Q(\bm{x},\bm{v})\right\rangle,$ (5)
where $\rho(\bm{x})=\left\langle f(\bm{x},\bm{v})\right\rangle$ denotes the macroscopic density function.
2.2 Random Feature Method
Before proposing the method in this paper, we will give a brief introduction to the RFM. Consider the following PDE in a bounded domain $\Omega\subset\mathbb{R}^{d}$:
$\mathcal{A}u(\bm{y})=0,\;\bm{y}\in\Omega,$ (6)
where $\mathcal{A}$ is a differential operator.
In general, the RFM represents the solution $u(\bm{y})$ by a two-layer neural networks with the inner parameters held fixed and chosen randomly. First, we introduce a hypercube $\Omega_{c}=\prod_{i=1}^{d}[a_{i},b_{i}]$ of proper size to completely enclose the domain $\Omega$. Here, the notation $\prod_{i=1}^{d}[a_{i},b_{i}]$ denotes the Cartesian product of $d$ intervals $[a_{i},b_{i}]$, that is, $[a_{1},b_{1}]\times[a_{2},b_{2}]\times\cdots\times[a_{d},b_{d}]$. Then we partition the hypercube $\Omega_{c}$ into $M$ non-overlapping hyper-rectangles $\Omega_{i}$:
$\Omega_{c}=\bigcup_{i=1}^{M}\Omega_{i},\;\Omega_{i}=\prod_{j=1}^{d}[a_{ij},b_{ij}],$ (7)
which is analogous to the finite element method and one can generalize the partition to a more general shape if necessary. Assume $\bm{y}=(y_{1},\cdots,y_{d})^{T}\in\Omega$, we apply the following transformation to obtain a partition dependent normalized vector $\tilde{\bm{y}}_{i}=(\tilde{y}_{i1},\cdots,\tilde{y}_{id})^{T}$:
$\tilde{y}_{ij}=2\frac{y_{j}-a_{ij}}{b_{ij}-a_{ij}}-1,\;i=1,\cdots,M,\;j=1,\cdots,d.$ (8)
Denote the center and radius of the hyper-rectangle $\Omega_{i}$ by:
$\bm{\mu}_{i}=\left(\frac{b_{i1}+a_{i1}}{2},\cdots,\frac{b_{id}+a_{id}}{2}\right)^{T},\;\bm{\sigma}_{i}=\left(\frac{b_{i1}-a_{i1}}{2},\cdots,\frac{b_{id}-a_{id}}{2}\right)^{T},$ (9)
then the transformation can be rewritten in the following vector form:
$\tilde{\bm{y}}_{i}=\frac{\bm{y}-\bm{\mu}_{i}}{\bm{\sigma}_{i}},\;i=1,\cdots,M.$ (10)
We can find that the above transformation maps the subdomain $\Omega_{i}$ to the hypercube $[-1,1]^{d}$.
The RFM consists of two parts: the PoU functions and the random feature functions. The PoU functions $\{\psi_{i}(\bm{y})\}_{i=1}^{M}$ is construct through the tensor product of univariate function:
$\psi_{i}(\bm{y})=\prod_{j=1}^{d}\varphi(\tilde{y}_{ij}),\;i=1,\cdots,M,$ (11)
where $\varphi$ is chosen in [39] as:
$\varphi_{a}(z)=\begin{cases}1,&|z|\leq 1,\\ 0,&\text{else},\end{cases}\;\text{or}\;\varphi_{b}(z)=\begin{cases}1,&|z|\leq\frac{3}{4},\\ \displaystyle\frac{1-\sin(2\pi|z|)}{2},&\frac{3}{4}\leq|z|\leq\frac{5}{4},\\ 0,&\text{else}.\end{cases}$ (12)
[…]
$\phi_{ij}(\bm{y})=\sigma\left(\bm{w}_{ij}\cdot\tilde{\bm{y}}_{i}+b_{ij}\right),\;i=1,\cdots,M,\;j=1,\cdots,J_{i},$ (13)
where $\bm{w}_{ij}\in\mathbb{R}^{d}$ and $b_{ij}\in\mathbb{R}$ are randomly generated with the uniform distribution in $[-B,B]$ and fixed parameters, and $\sigma$ is a scalar function, named activation function.
Finally, the RFM approximates the solution $u(\bm{y})$ through the linear combination of random feature functions together with PoU functions:
$u(\bm{y})\approx u_{M}(\bm{y})=\sum_{i=1}^{M}\psi_{i}(\bm{y})\sum_{j=1}^{J_{i}}u_{ij}\phi_{ij}(\bm{y}),$ (14)
[…]
$\min_{\theta}\|\mathcal{A}u_{M}(\bm{y})\|_{2}^{2},$ (15)
where $\|\cdot\|_{2}$ denotes the $\ell^{2}$ norm. Besides, the boundary and/or initial conditions which are denoted by $\mathcal{B}u_{M}(\bm{y})=0$ can be incorporated into the least squares problem (15) as constraints.
To solve the least squares problem (15), it is necessary to generate several collocation points for both $\Omega$ and $\partial\Omega$, respectively. Assume the interior collocation points are $\{\bm{y}_{\text{int}}^{k}\}_{1\leq k\leq N_{\text{int}}}$, and the boundary and/or initial collocation points are $\{\bm{y}_{\text{bdy}}^{k}\}_{1\leq k\leq N_{\text{bdy}}}$. Then the discrete least squares problem (15) can be written as:
$\min_{\theta}\sum_{k=1}^{N_{\text{int}}}\lambda_{\text{int}}^{k}|\mathcal{A}u_{M}(\bm{y}_{\text{int}}^{k})|^{2}+\sum_{k=1}^{N_{\text{bdy}}}\lambda_{\text{bdy}}^{k}|\mathcal{B}u_{M}(\bm{y}_{\text{bdy}}^{k})|^{2},$ (16)
where $\{\lambda_{\text{int}}^{k}\}$ and $\{\lambda_{\text{bdy}}^{k}\}$ are regularization parameters. The setting of the regularization parameters can be found in [39]. When the operators $\mathcal{A}$ and $\mathcal{B}$ are linear, the minimization problem (16) can be solved by the standard linear least squares method.
We present a schematic diagram of the RFM in Figure 1.
Figure 1: Schematic diagram of the random feature method.
2.3 Difficulty of RFM to resolve small scales
In this section, we demonstrate that the vanilla RFM has tremendous difficulty in resolving small scales in the multiscale radiative transfer equation through a simple example. Consider the following one-dimensional radiative transfer equation studied in [27]:
$\begin{cases}\displaystyle\varepsilon v\cdot\nabla_{x}f=\left\langle f\right\rangle-f-\varepsilon v,\;(x,v)\in[0,1]\times[-1,1],\\ f(0,v>0)=1,\;f(1,v<0)=0.\end{cases}$ (17)
The exact solution of the above equation is $f_{\text{ex}}(x,v)=1-x$. We apply the RFM to solve the radiative transfer equation (17), that is, we approximate the solution $f(x,v)$ by the following formula:
$f(x,v)\approx f_{M}(x,v)=\sum_{i=1}^{M}\psi_{i}(x,v)\sum_{j=1}^{J_{i}}f_{ij}\phi_{ij}(x,v).$ (18)
[…]
$\min_{\{f_{ij}\}_{1\leq i\leq M,\;1\leq j\leq J_{i}}}\mathcal{R}^{\varepsilon}:=\mathcal{R}_{\text{int}}^{\varepsilon}+\mathcal{R}_{\text{bdy}}^{\varepsilon},$ (19)
where $\mathcal{R}_{\text{int}}^{\varepsilon}$ and $\mathcal{R}_{\text{bdy}}^{\varepsilon}$ are defined as:
$\mathcal{R}_{\text{int}}^{\varepsilon}=\sum_{k=1}^{N_{\text{int}}}\lambda_{\text{int}}^{k}|\varepsilon v_{\text{int}}^{k}\cdot\nabla_{x}f_{M}(x_{\text{int}}^{k},v_{\text{int}}^{k})-\left\langle f_{M}\right\rangle(x_{\text{int}}^{k})+f_{M}(x_{\text{int}}^{k},v_{\text{int}}^{k})+\varepsilon v_{\text{int}}^{k}|^{2},$ (20)
and
$\mathcal{R}_{\text{bdy}}^{\varepsilon}=\sum_{k=1}^{N_{\text{bdy}}}\lambda_{\text{bdy}}^{k}|f_{M}(x_{\text{bdy}}^{k},v_{\text{bdy}}^{k})-f_{\text{bdy}}^{k}|^{2},$ (21)
respectively. Here, $\{(x_{\text{int}}^{k},v_{\text{int}}^{k})\}_{1\leq k\leq N_{\text{int}}}$ and $\{(x_{\text{bdy}}^{k},v_{\text{bdy}}^{k})\}_{1\leq k\leq N_{\text{bdy}}}$ are the interior and boundary collocation points, and $\{f_{\text{bdy}}^{k}\}_{1\leq k\leq N_{\text{bdy}}}$ are the boundary values. The regularization parameters $\{\lambda_{\text{int}}^{k}\}$ and $\{\lambda_{\text{bdy}}^{k}\}$ are positive constants.
Denote the linear operator
$\mathcal{T}^{\varepsilon}f(x,v)=\varepsilon v\cdot\nabla_{x}f(x,v)-\left\langle f\right\rangle(x)+f(x,v),$ (22)
we have
$\mathcal{T}^{\varepsilon}f_{M}=\mathcal{T}^{\varepsilon}\left(\sum_{i=1}^{M}\psi_{i}(x,v)\sum_{j=1}^{J_{i}}f_{ij}\phi_{ij}(x,v)\right)=\sum_{i=1}^{M}\sum_{j=1}^{J_{i}}f_{ij}\mathcal{T}^{\varepsilon}(\psi_{i}\phi_{ij}).$ (23)
Introduce the notation $s^{\text{int},k}_{ij}:=\mathcal{T}^{\varepsilon}(\psi_{i}\phi_{ij})(x_{\text{int}}^{k},v_{\text{int}}^{k})$ and $s^{\text{bdy},k}_{ij}:=\psi_{i}(x_{\text{bdy}}^{k},v_{\text{bdy}}^{k})\phi_{ij}(x_{\text{bdy}}^{k},v_{\text{bdy}}^{k})$, we can defined the matrix $\bm{A}$ and the vector $\bm{b}$
$\bm{A}=\begin{bmatrix}\bm{A}_{\text{int}}\\ \bm{A}_{\text{bdy}}\end{bmatrix}\in\mathbb{R}^{N\times Q},\;\bm{b}=\begin{bmatrix}\bm{b}_{\text{int}}\\ \bm{b}_{\text{bdy}}\end{bmatrix}\in\mathbb{R}^{N},\;N=N_{\text{int}}+N_{\text{bdy}},$ (24)
with
$\bm{A}_{\text{int}}=\begin{bmatrix}s^{\text{int},1}_{11}&\cdots&s^{\text{int},1}_{1J_{1}}&s^{1}_{21}&\cdots&s^{\text{int},1}_{MJ_{M}}\\ s^{\text{int},2}_{11}&\cdots&s^{\text{int},2}_{1J_{1}}&s^{2}_{21}&\cdots&s^{\text{int},2}_{MJ_{M}}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ s^{\text{int},N_{\text{int}}}_{11}&\cdots&s^{\text{int},N_{\text{int}}}_{1J_{1}}&s^{\text{int},N_{\text{int}}}_{21}&\cdots&s^{\text{int},N_{\text{int}}}_{MJ_{M}}\end{bmatrix},\,\bm{b}_{\text{int}}=-\varepsilon\begin{bmatrix}v_{\text{int}}^{1}\\ v_{\text{int}}^{2}\\ \vdots\\ v_{\text{int}}^{N_{\text{int}}}\end{bmatrix},$ (25)
$\bm{A}_{\text{bdy}}=\begin{bmatrix}s^{\text{bdy},1}_{11}&\cdots&s^{\text{bdy},1}_{1J_{1}}&s^{\text{bdy},1}_{21}&\cdots&s^{\text{bdy},1}_{MJ_{M}}\\ s^{\text{bdy},2}_{11}&\cdots&s^{\text{bdy},2}_{1J_{1}}&s^{\text{bdy},2}_{21}&\cdots&s^{\text{bdy},2}_{MJ_{M}}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ s^{\text{bdy},N_{\text{bdy}}}_{11}&\cdots&s^{\text{bdy},N_{\text{bdy}}}_{1J_{1}}&s^{\text{bdy},N_{\text{bdy}}}_{21}&\cdots&s^{\text{bdy},N_{\text{bdy}}}_{MJ_{M}}\end{bmatrix},\;\bm{b}_{\text{bdy}}=\begin{bmatrix}f_{\text{bdy}}^{1}\\ f_{\text{bdy}}^{2}\\ \vdots\\ f_{\text{bdy}}^{N_{\text{bdy}}}\end{bmatrix}.$ (26)
Let the vector of coefficients be
$\bm{f}=\begin{bmatrix}f_{11}&\cdots&f_{1J_{1}}&f_{21}&\cdots&f_{MJ_{M}}\end{bmatrix}^{T},$ (27)
the optimal coefficients $\bm{f}$ of the problem (19) can be obtained by
$\bm{f}^{*}=\min_{\bm{f}}\|\bm{A}\bm{f}-\bm{b}\|_{2}^{2}.$ (28)
The solution $\bm{f}^{*}$ can be obtained by the standard linear least squares method.
First, we consider the problem (17) with $\varepsilon=1$ to validate the effectiveness of the RFM. Here, we set the number of partitions $M=1$ with the PoU function $\varphi_{a}$. We choose the $\tanh$ function as our activation function and the number of random feature functions $J_{i}=2^{n},n=4,5,6,7$. The weights $\bm{w}_{ij}$ and biases $b_{ij}$ are randomly generated with $B=1$. Besides, the collocation points are uniform grids with $(N_{x},N_{v})=(32,64)$. The integration of the operator $\mathcal{L}$ is approximated by Gauss-Legendre quadrature with $16$ points.
The relative $\ell^{2}$ error of the solution $f_{M}$ concerning the degrees of freedom $Q$ is shown in Figure 2. One can observe that the error decreases exponentially to the degrees of freedom $Q$. We also plot the solution $f_{M}$ obtained by the random feature method with $J_{i}=2^{7}$ and reference solution $f_{\text{ex}}$ in Figure 3. The relative $\ell^{2}$ error of the RFM is $3.99\times 10^{-7}$.
Figure 2: Relative $\ell^{2}$ error of the RFM solution $f_{M}$ with respect to the degrees of freedom $Q$.
Figure 3: Reference solution v.s. RFM solution.
Next, we test a sequential of problems (17) with $\varepsilon=10^{-2^{i}}$, $i=1,2,3,4$ with the same setting as the previous problem except for the number of random features $J_{i}$ and the collocation points $(N_{x},N_{v})$. We record the relative $\ell^{2}$ error of the RFM solution to the degrees of freedom and the number of collocation points in Table 1 and Table 2, respectively. One can find that the small-scale problems ($\varepsilon\ll 1$) are much more difficult to resolve than the large-scale problems in terms of the degrees of freedom and the number of collocation points.
Table 1: The dependency of the RFM on degrees of freedom. The collocation points are uniform grids with $(N_{x},N_{v})=(128,256)$.
$\varepsilon$ $J_{i}$
$16$ $32$ $64$ $128$ $256$
[…]
Table 2: The dependency of the RFM on the number of collocation points. The number of random feature functions $J_{i}=128$.
$\varepsilon$ $(N_{x},\,N_{v})$
(16, 32)1632(16,\,32)( 16 , 32 ) (32, 64)3264(32,\,64)( 32 , 64 ) (64, 128)64128(64,\,128)( 64 , 128 ) (128, 256)128256(128,\,256)( 128 , 256 )
[…]
Consider the problem (17) with small scale, i.e., $\varepsilon\ll 1$. It can be observed that any function $f$ independent of $v$, which satisfies the inflow boundary condition, such as $f=(1-x)^{n}$ for $n\geq 2$, will lead to the following:
$\|\varepsilon v\cdot\nabla_{x}f-\left\langle f\right\rangle+f+\varepsilon v\|_{2}^{2}=\mathcal{O}(\varepsilon^{2}),$ (29)
but
$\|f-f_{\text{ex}}\|_{2}^{2}=\mathcal{O}(1).$ (30)
To resolve the small-scale information $\mathcal{O}(\varepsilon^{2})$, an increase in both the degrees of freedom and the number of collocation points is required. This leads to a large matrix $\bm{A}$ with a high condition number. The high condition number of the matrix $\bm{A}$ will make the least squares problem (28) ill-conditioned, and the solution $\bm{f}^{*}$ may not be accurate enough. The above analysis implies that the vanilla RFM based on least squares formulation for the radiative transfer equation (1) has tremendous difficulty in resolving small scales in the radiative transfer equation. This is not only due to the ill-conditioning of the least squares problem, but also the lack of capturing the small-scale part of the solution $f(x,v)$. Such a similar phenomenon is also observed in the numerical experiments in [28, 27] when solving the multiscale kinetic equations with small parameters through deep neural networks.
3 Asymptotic-Preserving Random Feature Method
To address the issue discussed in the previous section, we propose a new random feature method based on a micro-macro decomposition [45, 46], called the asymptotic-preserving random feature method. We introduce the definition of the asymptotic-preserving random feature method as follows (see Figure 4):
Definition 1.
Assume $\mathcal{F^{\varepsilon}}$ is the microscopic equation that depends on the scale parameter $\varepsilon$ and $\mathcal{F}^{0}$ is its macroscopic limit equation as $\varepsilon\to 0$. The latent solution of $\mathcal{F^{\varepsilon}}$ is approximated by the RFM and solved by the least squares minimization problem $\mathcal{R}(\mathcal{F^{\varepsilon}})$. The asymptotic limit of $\mathcal{R}(\mathcal{F^{\varepsilon}})$ as $\varepsilon\to 0$, if exists, is denoted by $\mathcal{R}(\mathcal{F}^{0})$. If $\mathcal{R}(\mathcal{F}^{0})$ is a “good” least squares formulation of $\mathcal{F}^{0}$, which can capture the multiscale information, then the method is called asymptotic-preserving.
Figure 4: Schematic diagram of the asymptotic-preserving random feature method.
Let us revisit the radiative transfer equation (1),
$\bm{v}\cdot\nabla_{\bm{x}}f(\bm{x},\bm{v})=\frac{\sigma_{s}(\bm{x})}{\varepsilon}\mathcal{L}f(\bm{x},\bm{v})-\varepsilon\sigma_{a}(\bm{x})f(\bm{x},\bm{v})+\varepsilon Q(\bm{x},\bm{v}),$ (31)
and we decompose the distribution function $f(\bm{x},\bm{v})$ into two parts:
$f(\bm{x},\bm{v})=\rho(\bm{x})+\varepsilon g(\bm{x},\bm{v}),$ (32)
where $\rho(\bm{x})=\left\langle f(\bm{x},\bm{v})\right\rangle$ denotes the equilibrium part, and $g(\bm{x},\bm{v})$ denotes the non-equilibrium part satisfying $\left\langle g(\bm{x},\bm{v})\right\rangle=0$.
Substitute the decomposition (32) into the radiative transfer equation (31), and according to the properties of the operator $\mathcal{L}$, we obtain the following equations:
$\bm{v}\cdot\nabla_{\bm{x}}(\rho(\bm{x})+\varepsilon g(\bm{x},\bm{v}))=\sigma_{s}(\bm{x})\mathcal{L}g(\bm{x},\bm{v})-\varepsilon\sigma_{a}(\bm{x})(\rho(\bm{x})+\varepsilon g(\bm{x},\bm{v}))+\varepsilon Q(\bm{x},\bm{v}).$ (33)
Integrating the above equation over the velocity space $\mathbb{S}^{d-1}$, we have
$\left\langle\bm{v}\cdot\nabla_{\bm{x}}g\right\rangle=-\sigma_{a}\rho+\left\langle Q\right\rangle.$ (34)
Define the orthogonal projection operator
$\Pi:\Pi(\cdot)(\bm{v})=\left\langle\cdot\right\rangle,$ (35)
and the identity operator Id, one can apply the operator $\text{Id}-\Pi$ to the equation (33) to obtain the following equation:
$\bm{v}\cdot\nabla_{\bm{x}}\rho+\varepsilon(\text{Id}-\Pi)(\bm{v}\cdot\nabla_{\bm{x}}g)=\sigma_{s}\mathcal{L}g-\varepsilon^{2}\sigma_{a}g+\varepsilon(Q-\left\langle Q\right\rangle).$ (36)
Therefore, equation (34) and (36) constitute the micro-macro system of the radiative transfer equation (31):
$\begin{cases}\left\langle\bm{v}\cdot\nabla_{\bm{x}}g\right\rangle+\sigma_{a}\rho=\left\langle Q\right\rangle,\\ \bm{v}\cdot\nabla_{\bm{x}}\rho+\varepsilon(\text{Id}-\Pi)(\bm{v}\cdot\nabla_{\bm{x}}g)=\sigma_{s}\mathcal{L}g-\varepsilon^{2}\sigma_{a}g+\varepsilon(Q-\left\langle Q\right\rangle).\end{cases}$ (37)
Our idea is to approximate $\rho(\bm{x})$ and $g(\bm{x},\bm{v})$ by the following formula:
$\rho(\bm{x})\approx\rho_{M}(\bm{x})=\sum_{i=1}^{M^{\rho}}\psi_{i}^{\rho}(\bm{x})\sum_{j=1}^{J_{i}^{\rho}}\rho_{ij}\phi_{ij}^{\rho}(\bm{x}),$ (38)
and
$g(\bm{x},\bm{v})\approx g_{M}(\bm{x},\bm{v})=\sum_{i=1}^{M^{g}}\psi_{i}^{g}(\bm{x},\bm{v})\sum_{j=1}^{J_{i}^{g}}g_{ij}\phi_{ij}^{g}(\bm{x},\bm{v}),$ (39)
[…]
$\min_{\bm{\theta}}\mathcal{R}^{\varepsilon}:=\mathcal{R}_{\text{int}}^{\varepsilon}+\mathcal{R}_{\text{bdy}}^{\varepsilon},$ (40)
where $\mathcal{R}_{\text{int}}^{\varepsilon}$ and $\mathcal{R}_{\text{bdy}}^{\varepsilon}$ are defined as:
$\mathcal{R}_{\text{int}}^{\varepsilon}$ $=\sum_{k=1}^{N_{\text{int}}}\lambda_{\text{int},1}^{k}|\left\langle\bm{v}\cdot\nabla_{\bm{x}}g_{M}\right\rangle(\bm{x}_{\text{int}}^{k})+\sigma_{a}(\bm{x}_{\text{int}}^{k})\rho_{M}(\bm{x}_{\text{int}}^{k})-\left\langle Q\right\rangle(\bm{x}_{\text{int}}^{k})|^{2}$ (41)
$+\sum_{k=1}^{N_{\text{int}}}\lambda_{\text{int},2}^{k}|\bm{v}_{\text{int}}^{k}\cdot\nabla_{\bm{x}}\rho_{M}(\bm{x}_{\text{int}}^{k})+\varepsilon(\text{Id}-\Pi)(\bm{v}_{\text{int}}^{k}\cdot\nabla_{\bm{x}}g_{M}(\bm{x}_{\text{int}}^{k},\bm{v}_{\text{int}}^{k}))$
$\qquad\qquad-\sigma_{s}(\bm{x}_{\text{int}}^{k})\mathcal{L}g_{M}(\bm{x}_{\text{int}}^{k},\bm{v}_{\text{int}}^{k})+\varepsilon^{2}\sigma_{a}(\bm{x}_{\text{int}}^{k})g_{M}(\bm{x}_{\text{int}}^{k},\bm{v}_{\text{int}}^{k})$
$\qquad\qquad-\varepsilon(Q(\bm{x}_{\text{int}}^{k},\bm{v}_{\text{int}}^{k})-\left\langle Q\right\rangle(\bm{x}_{\text{int}}^{k}))|^{2},$
and
$\mathcal{R}_{\text{bdy}}^{\varepsilon}=\sum_{k=1}^{N_{\text{bdy}}}\lambda_{\text{bdy}}^{k}|\rho_{M}(\bm{x}_{\text{bdy}}^{k})+\varepsilon g_{M}(\bm{x}_{\text{bdy}}^{k},\bm{v}_{\text{bdy}}^{k})-f_{\text{bdy}}^{k}|^{2},$ (42)
respectively. Here, $\{(\bm{x}_{\text{int}}^{k},\bm{v}_{\text{int}}^{k})\}_{1\leq k\leq N_{\text{int}}}$ and $\{(\bm{x}_{\text{bdy}}^{k},\bm{v}_{\text{bdy}}^{k})\}_{1\leq k\leq N_{\text{bdy}}}$ are the interior and boundary collocation points, and $\{f_{\text{bdy}}^{k}\}_{1\leq k\leq N_{\text{bdy}}}$ are the boundary values. $\{\lambda_{\text{int},1}^{k}\},\{\lambda_{\text{int},2}^{k}\}$ and $\{\lambda_{\text{bdy}}^{k}\}$ are regularization parameters. The degrees of freedom of the APRFM are $Q=\sum_{i=1}^{M^{\rho}}J_{i}^{\rho}+\sum_{i=1}^{M^{g}}J_{i}^{g}$.
Denote the linear operators
$\mathcal{T}_{11}^{\varepsilon}\rho=\sigma_{a}(\bm{x})\rho(\bm{x}),$ (43)
$\mathcal{T}_{12}^{\varepsilon}g=\left\langle\bm{v}\cdot\nabla_{\bm{x}}g(\bm{x},\bm{v})\right\rangle,$ (44)
$\mathcal{T}_{21}^{\varepsilon}\rho=\bm{v}\cdot\nabla_{\bm{x}}\rho(\bm{x}),$ (45)
and
$\mathcal{T}_{22}^{\varepsilon}g=\varepsilon(\text{Id}-\Pi)(\bm{v}\cdot\nabla_{\bm{x}}g(\bm{x},\bm{v}))-\sigma_{s}(\bm{x})\mathcal{L}g(\bm{x},\bm{v})+\varepsilon^{2}\sigma_{a}(\bm{x})g(\bm{x},\bm{v}).$ (46)
We have
$\mathcal{T}_{11}^{\varepsilon}\rho_{M}=\sum_{i=1}^{M^{\rho}}\sum_{j=1}^{J_{i}^{\rho}}\rho_{ij}\mathcal{T}_{11}^{\varepsilon}(\psi_{i}^{\rho}\phi_{ij}^{\rho}),\;\mathcal{T}_{12}^{\varepsilon}g_{M}=\sum_{i=1}^{M^{g}}\sum_{j=1}^{J_{i}^{g}}g_{ij}\mathcal{T}_{12}^{\varepsilon}(\psi_{i}^{g}\phi_{ij}^{g}),$ (47)
and
$\mathcal{T}_{21}^{\varepsilon}\rho_{M}=\sum_{i=1}^{M^{\rho}}\sum_{j=1}^{J_{i}^{\rho}}\rho_{ij}\mathcal{T}_{21}^{\varepsilon}(\psi_{i}^{\rho}\phi_{ij}^{\rho}),\;\mathcal{T}_{22}^{\varepsilon}g_{M}=\sum_{i=1}^{M^{g}}\sum_{j=1}^{J_{i}^{g}}g_{ij}\mathcal{T}_{22}^{\varepsilon}(\psi_{i}^{g}\phi_{ij}^{g}).$ (48)
Introduce the notations
$^{11}s^{\text{int},k}_{ij}:=\mathcal{T}_{11}^{\varepsilon}(\psi_{i}^{\rho}\phi_{ij}^{\rho})(\bm{x}_{\text{int}}^{k},\bm{v}_{\text{int}}^{k}),$ (49)
$^{12}s^{\text{int},k}_{ij}:=\mathcal{T}_{12}^{\varepsilon}(\psi_{i}^{g}\phi_{ij}^{g})(\bm{x}_{\text{int}}^{k},\bm{v}_{\text{int}}^{k}),$ (50)
$^{21}s^{\text{int},k}_{ij}:=\mathcal{T}_{21}^{\varepsilon}(\psi_{i}^{\rho}\phi_{ij}^{\rho})(\bm{x}_{\text{int}}^{k},\bm{v}_{\text{int}}^{k}),$ (51)
$^{22}s^{\text{int},k}_{ij}:=\mathcal{T}_{22}^{\varepsilon}(\psi_{i}^{g}\phi_{ij}^{g})(\bm{x}_{\text{int}}^{k},\bm{v}_{\text{int}}^{k}),$ (52)
and
$^{31}s^{\text{bdy},k}_{ij}:=(\psi_{i}^{\rho}\phi_{ij}^{\rho})(\bm{x}_{\text{bdy}}^{k}),$ (53)
$^{32}s^{\text{bdy},k}_{ij}:=(\varepsilon\psi_{i}^{g}\phi_{ij}^{g})(\bm{x}_{\text{bdy}}^{k},\bm{v}_{\text{bdy}}^{k}),$ (54)
then we construct the matrix $\bm{A}$ and the vector $\bm{b}$ as follows:
$\bm{A}=\begin{bmatrix}\bm{A}_{\text{int}}\\ \bm{A}_{\text{bdy}}\end{bmatrix}\in\mathbb{R}^{N\times Q},\;\bm{b}=\begin{bmatrix}\bm{b}_{\text{int}}\\ \bm{b}_{\text{bdy}}\end{bmatrix}\in\mathbb{R}^{N},\;N=2N_{\text{int}}+N_{\text{bdy}},$ (55)
$\bm{A}_{\text{int}}=\begin{bmatrix}{}^{11}\bm{A}_{\text{int}}^{1}&{}^{12}\bm{A}_{\text{int}}^{1}\\ {}^{21}\bm{A}_{\text{int}}^{1}&{}^{22}\bm{A}_{\text{int}}^{1}\\ {}^{11}\bm{A}_{\text{int}}^{2}&{}^{12}\bm{A}_{\text{int}}^{2}\\ {}^{21}\bm{A}_{\text{int}}^{2}&{}^{22}\bm{A}_{\text{int}}^{2}\\ \vdots&\vdots\\ {}^{11}\bm{A}_{\text{int}}^{N_{\text{int}}}&{}^{12}\bm{A}_{\text{int}}^{N_{\text{int}}}\\ {}^{21}\bm{A}_{\text{int}}^{N_{\text{int}}}&{}^{22}\bm{A}_{\text{int}}^{N_{\text{int}}}\end{bmatrix},\;\bm{A}_{\text{bdy}}=\begin{bmatrix}{}^{31}\bm{A}_{\text{bdy}}^{1}&{}^{32}\bm{A}_{\text{bdy}}^{1}\\ {}^{31}\bm{A}_{\text{bdy}}^{2}&{}^{32}\bm{A}_{\text{bdy}}^{2}\\ \vdots&\vdots\\ {}^{31}\bm{A}_{\text{bdy}}^{N_{\text{bdy}}}&{}^{32}\bm{A}_{\text{bdy}}^{N_{\text{bdy}}}\end{bmatrix},$ (56)
and
$\bm{b}_{\text{int}}=\begin{bmatrix}\left\langle Q\right\rangle(\bm{x}_{\text{int}}^{1})\\ \varepsilon(Q(\bm{x}_{\text{int}}^{1},\bm{v}_{\text{int}}^{1})-\left\langle Q\right\rangle(\bm{x}_{\text{int}}^{1}))\\ \left\langle Q\right\rangle(\bm{x}_{\text{int}}^{2})\\ \varepsilon(Q(\bm{x}_{\text{int}}^{2},\bm{v}_{\text{int}}^{2})-\left\langle Q\right\rangle(\bm{x}_{\text{int}}^{2}))\\ \vdots\\ \left\langle Q\right\rangle(\bm{x}_{\text{int}}^{N_{\text{int}}})\\ \varepsilon(Q(\bm{x}_{\text{int}}^{N_{\text{int}}},\bm{v}_{\text{int}}^{N_{\text{int}}})-\left\langle Q\right\rangle(\bm{x}_{\text{int}}^{N_{\text{int}}}))\end{bmatrix},\;\bm{b}_{\text{bdy}}=\begin{bmatrix}f_{\text{bdy}}^{1}\\ f_{\text{bdy}}^{2}\\ \vdots\\ f_{\text{bdy}}^{N_{\text{bdy}}}\end{bmatrix},$ (57)
where the vectors ${}^{11}\bm{A}_{\text{int}}^{k},^{12}\bm{A}_{\text{int}}^{k},^{21}\bm{A}_{\text{int}}^{k},^{22}\bm{A}_{\text{int}}^{k},^{31}\bm{A}_{\text{bdy}}^{k},^{32}\bm{A}_{\text{bdy}}^{k}$ are defined as
${}^{11}\bm{A}_{\text{int}}^{k}=\begin{bmatrix}{}^{11}s^{\text{int},k}_{11}&\cdots&{}^{11}s^{\text{int},k}_{1J_{1}^{\rho}}&{}^{11}s^{\text{int},k}_{21}&\cdots&{}^{11}s^{\text{int},k}_{M^{\rho}J_{M^{\rho}}}\end{bmatrix},$ (58)
${}^{12}\bm{A}_{\text{int}}^{k}=\begin{bmatrix}{}^{12}s^{\text{int},k}_{11}&\cdots&{}^{12}s^{\text{int},k}_{1J_{1}^{g}}&{}^{12}s^{\text{int},k}_{21}&\cdots&{}^{12}s^{\text{int},k}_{M^{g}J_{M^{g}}}\end{bmatrix},$ (59)
${}^{21}\bm{A}_{\text{int}}^{k}=\begin{bmatrix}{}^{21}s^{\text{int},k}_{11}&\cdots&{}^{21}s^{\text{int},k}_{1J_{1}^{\rho}}&{}^{21}s^{\text{int},k}_{21}&\cdots&{}^{21}s^{\text{int},k}_{M^{\rho}J_{M^{\rho}}}\end{bmatrix},$ (60)
${}^{22}\bm{A}_{\text{int}}^{k}=\begin{bmatrix}{}^{22}s^{\text{int},k}_{11}&\cdots&{}^{22}s^{\text{int},k}_{1J_{1}^{g}}&{}^{22}s^{\text{int},k}_{21}&\cdots&{}^{22}s^{\text{int},k}_{M^{g}J_{M^{g}}}\end{bmatrix},$ (61)
${}^{31}\bm{A}_{\text{bdy}}^{k}=\begin{bmatrix}{}^{31}s^{\text{bdy},k}_{11}&\cdots&{}^{31}s^{\text{bdy},k}_{1J_{1}^{\rho}}&{}^{31}s^{\text{bdy},k}_{21}&\cdots&{}^{31}s^{\text{bdy},k}_{M^{\rho}J_{M^{\rho}}}\end{bmatrix},$ (62)
${}^{32}\bm{A}_{\text{bdy}}^{k}=\begin{bmatrix}{}^{32}s^{\text{bdy},k}_{11}&\cdots&{}^{32}s^{\text{bdy},k}_{1J_{1}^{g}}&{}^{32}s^{\text{bdy},k}_{21}&\cdots&{}^{32}s^{\text{bdy},k}_{M^{g}J_{M^{g}}}\end{bmatrix}.$ (63)
Let $\bm{\theta}=[\bm{\rho},\bm{g}]^{T}$, with
$\bm{\rho}=\begin{bmatrix}\rho_{11}&\cdots&\rho_{1J_{1}^{\rho}}&\rho_{21}&\cdots&\rho_{M^{\rho}J_{M^{\rho}}}\end{bmatrix},$ (64)
and
$\bm{g}=\begin{bmatrix}g_{11}&\cdots&g_{1J_{1}^{g}}&g_{21}&\cdots&g_{M^{g}J_{M^{g}}}\end{bmatrix},$ (65)
then the least squares problem (40) can be written as
$\bm{\theta}^{*}=\min_{\bm{\theta}}\|\bm{A}\bm{\theta}-\bm{b}\|_{2}^{2}.$ (66)
The algorithm of the APRFM is summarized in Algorithm 1.
Input: The set of collocation points $\{(\bm{x}_{\text{int}}^{k},\bm{v}_{\text{int}}^{k})\}_{1\leq k\leq N_{\text{int}}}$, $\{(\bm{x}_{\text{bdy}}^{k},\bm{v}_{\text{bdy}}^{k})\}_{1\leq k\leq N_{\text{bdy}}}$, $\{f_{\text{bdy}}^{k}\}_{1\leq k\leq N_{\text{bdy}}}$; The number of PoU functions $M^{\rho}$, $M^{g}$ and random feature functions $\{J_{i}^{\rho}\}_{1\leq i\leq M^{\rho}}$, $\{J_{i}^{g}\}_{1\leq i\leq M^{g}}$; The range of uniform distribution $[-B,B]$ for the initialization of neural network parameters; The Gauss-Legendre quadrature points and weights $\{(\omega_{j},\bm{v}_{j})\}_{j=1}^{N_{q}}$.
Output: The coefficients $\bm{\theta}^{*}$.
1
2Initialize the weights and biases of the neural networks $\{\phi_{ij}^{\rho}\},\{\phi_{ij}^{g}\}$ randomly according to the uniform distribution $[-B,B]$ and keep them fixed;
3Set $k_{1}=1,k_{2}=1$;
4while $k_{1}\leq N_{\text{int}}$ do
5       for $i=1,\cdots,M^{\rho}$ do
6             for $j=1,\cdots,J_{i}^{\rho}$ do
7                   Compute ${}^{11}s^{\text{int},k_{1}}_{ij},^{21}s^{\text{int},k_{1}}_{ij}$ according to (49) and (51);
8             end for
9
10       end for
11      for $i=1,\cdots,M^{g}$ do
12             for $j=1,\cdots,J_{i}^{g}$ do
13                   Compute ${}^{12}s^{\text{int},k_{1}}_{ij},^{22}s^{\text{int},k_{1}}_{ij}$ according to (50) and (52);
14             end for
15
16       end for
17
18      $k_{1}=k_{1}+1$;
19 end while
20
21while $k_{2}\leq N_{\text{bdy}}$ do
22       for $i=1,\cdots,M^{\rho}$ do
23             for $j=1,\cdots,J_{i}^{\rho}$ do
24                   Compute ${}^{31}s^{\text{bdy},k_{2}}_{ij}$ according to (53);
25             end for
26
27       end for
28      for $i=1,\cdots,M^{g}$ do
29             for $j=1,\cdots,J_{i}^{g}$ do
30                   Compute ${}^{32}s^{\text{bdy},k_{2}}_{ij}$ according to (54);
31             end for
32
33       end for
34
35      $k_{2}=k_{2}+1$;
36 end while
37
38Construct the matrix $\bm{A}$ and the vector $\bm{b}$ according to (56) and (57);
39Solve the least squares problem (66) to obtain $\bm{\theta}^{*}$.
return $\bm{\theta}^{*}$.
Algorithm 1 APRFM based on Micro-Macro Decomposition.
4 Numerical results
In this section, we present the numerical results of the APRFM for the radiative transfer equation. To enhance the performance of APRFM, in this paper, we introduce a new 1D PoU function $\varphi$ as follows:
$\varphi(z)=\exp(-z^{2}),$ (67)
it can be observed that $\varphi$ is a $C^{\infty}$ function with no compact support. This PoU function is employed in the following numerical examples, except for Example 1. For all subsequent examples, unless explicitly mentioned otherwise, the default activation function is set as the $\tanh$ function. The collocation points are selected as uniform grids, denoted by $(N_{x},N_{v})$ for 1D cases and $(N_{x_{1}},N_{x_{2}},N_{v})$ for 2D cases. The domain partitioning is represented as $(M_{x},M_{v})$ for 1D cases and $(M_{x_{1}},M_{x_{2}},M_{v})$ for 2D cases. All reference solutions in the examples are computed using the finite difference method.
4.1 One-dimensional problems
Example 1
Let us consider the above example (17) in section 2.3:
$\begin{cases}\displaystyle\varepsilon v\cdot\nabla_{x}f=\left\langle f\right\rangle-f-\varepsilon v,\;(x,v)\in[0,1]\times[-1,1],\\ f(0,v>0)=1,\;f(1,v<0)=0,\end{cases}$ (68)
with the exact solution $f_{\text{ex}}=1-x$. For the kinetic regime ($\varepsilon=1$), we set $J_{i}^{\rho}=J_{i}^{g}=2^{n-1},n=4,5,6,7$ and $B=2.5$. Other settings are the same as Figure 2 in the previous section. The relative $\ell^{2}$ error of the APRFM solution to the degrees of freedom $Q=2^{n}$ is shown in Figure 5.
Figure 5: Relative $\ell^{2}$ error of the APRFM solution with respect to the degrees of freedom $Q$.
To investigate the dependency of the scale parameter in APRFM on degrees of freedom and collocation points, we consider $\varepsilon=10^{-2^{i}}$, $i=1,2,3,4$ with different random features $J_{i}^{\rho}=J_{i}^{g}=J$ and the collocation points $(N_{x},N_{v})$. We record the relative $\ell^{2}$ error of the APRFM solution with respect to the degrees of freedom and the number of collocation points in Table 3 and Table 4, respectively. We can observe that our APRFM is robust with respect to the scale parameter $\varepsilon$ and needs fewer degrees of freedom and collocation points to achieve high accuracy compared with the RFM in Table 1 and Table 2.
Table 3: The dependency of the APRFM on degrees of freedom. The collocation points are uniform grids with $(N_{x},N_{v})=(128,256)$.
$\varepsilon$ $J$
$8$ $16$ $32$ $64$ $128$
[…]
Table 4: The dependency of the APRFM on the number of collocation points. The number of random features $J=128$.
$\varepsilon$ $(N_{x},\,N_{v})$
(16, 32)1632(16,\,32)( 16 , 32 ) (32, 64)3264(32,\,64)( 32 , 64 ) (64, 128)64128(64,\,128)( 64 , 128 ) (128, 256)128256(128,\,256)( 128 , 256 )
[…]
Example 2
In this example, we consider the kinetic regime ($\varepsilon=1$) and intermediate regime ($\varepsilon=5\times 10^{-1}$) in 1D slab geometry without the presence of a source term:
$\begin{cases}\displaystyle\varepsilon v\cdot\nabla_{x}f=\left\langle f\right\rangle-f,\;(x,v)\in[0,1]\times[-1,1],\\ f(0,v>0)=1,\;f(1,v<0)=0.\end{cases}$ (69)
We plot the solution obtained by our APRFM and reference solution in Figure 6. We set $J_{i}^{\rho}=64,J_{i}^{g}=128$ and $B=1.5$. The number of interior collocation points is $(N_{x},N_{v})=(256,128)$, while the number of boundary collocation points is $N_{v}=64$. Additionally, we use $24$ points in the Gauss-Legendre quadrature. The domain is partitioned as $(M_{x},M_{v})=(2,4)$, i.e., $M^{\rho}=2,M^{g}=8$. The relative $\ell^{2}$ error of our APRFM is $7.87\times 10^{-3}$ for $\varepsilon=1$ and $5.70\times 10^{-3}$ for $\varepsilon=5\times 10^{-1}$, respectively. It can be observed that the errors of the APRFM are predominantly concentrated at $x=0,1$ and $v=0$.
(a) Reference solution v.s. APRFM solution ($\varepsilon=1$).
(b) Reference solution v.s. APRFM solution ($\varepsilon=5\times 10^{-1}$).
Figure 6: Plots of $f(x,v)$ in $(x,v)\in[0,1]\times[-1,1]$ by APRFM and finite difference method.
In the kinetic case ($\varepsilon=1$), we also compared our proposed APRFM with the APNN method [28, 27]. As shown in Figure 7, our approach not only achieves slightly lower relative $\ell^{2}$ error compared to the APNN method, but more importantly, it significantly outperforms in terms of computational efficiency, being an order of magnitude faster and requiring considerably fewer parameters. Owing to the uniform accuracy of our APRFM, the comparative results hold similarly for cases where $\varepsilon\ll 1$. This comparative experiment was conducted on a machine equipped with an NVIDIA A800 80G GPU and a 48-core Intel Xeon Gold 6342 CPU. The neural networks for $\rho$ and $g$ in the APNN method are 4-layer fully connected architectures, each with 128 units per hidden layer, utilizing the $tanh$ activation function. We trained the APNN method for 50000 iterations using the Adam optimizer, with a batch size of 8196 for interior points and 2048 for boundary points during the training process.
Figure 7: The comparison of APRFM and APNN.
Example 3
The following example involves a mixed-scale RTE:
$\begin{cases}{}\displaystyle v\cdot\nabla_{x}f=\frac{1}{\varepsilon(x)}\left(\left\langle f\right\rangle-f\right),\;(x,v)\in[0,1]\times[-1,1],\\ f(0,v>0)=0.5,\;f(1,v<0)=0.\end{cases}$ (70)
This case can verify the adaptability of our APRFM under mixed multi-scale conditions. Here, $\varepsilon(x)$ is a function that depends on the spatial variable and smoothly transitions from $\mathcal{O}(10^{-3})$ to $\mathcal{O}(1)$, defined as follows:
$\varepsilon(x)=10^{-2}+\frac{1}{2}\left[\tanh(6.5-11x)+\tanh(11x-4.5)\right].$ (71)
The plot of the function $\varepsilon(x)$ is presented in Figure 8.
Figure 8: Plot of $\varepsilon(x)$.
We plot the solution obtained by our APRFM and reference solution in Figure 9. We set $J_{i}^{\rho}=96,J_{i}^{g}=128$ and $B=2.5$. The number of interior collocation points is $(N_{x},N_{v})=(512,256)$, while the number of boundary collocation points is $N_{v}=64$. Additionally, we use $32$ points in the Gauss-Legendre quadrature. The domain is partitioned as $(M_{x},M_{v})=(2,4)$, i.e., $M^{\rho}=2,M^{g}=8$. The relative $\ell^{2}$ error of our APRFM is $2.88\times 10^{-2}$.
Figure 9: Reference solution v.s. APRFM solution.
4.2 Two-dimensional problems
Example 4
In this subsection, we consider a 2D square domain with $\bm{x}=(x_{1},x_{2})\in[-1,1]^{2},\bm{v}=(\cos(\alpha),\sin(\alpha))$.
$\begin{cases}{}\displaystyle\varepsilon\bm{v}\cdot\nabla_{\bm{x}}f(\bm{x},\bm{v})=\frac{1}{2\pi}\int_{|\bm{v}|=1}f(\bm{x},\bm{v}^{\prime})\mathop{}\!\mathrm{d}{\bm{v}^{\prime}}-f+\varepsilon^{2}G(\bm{x},\bm{v}),\\ f(-1,x_{2},\alpha)=e^{1-x_{2}},\;\alpha\in[0,\pi/2]\cup[3\pi/2,2\pi],\\ f(1,x_{2},\alpha)=e^{-1-x_{2}},\;\alpha\in[\pi/2,3\pi/2],\\ f(x_{1},-1,\alpha)=e^{-x_{1}+1},\;\alpha\in[0,\pi],\\ f(x_{1},1,\alpha)=e^{-x_{1}-1},\;\alpha\in[\pi,2\pi].\end{cases}$ (72)
Here, the source function is $G(\bm{x},\bm{v})=(-\cos(\alpha)-\sin(\alpha))\exp(-x-y)/\varepsilon$. In this example, the exact solution is given by $f_{\text{ex}}=\exp(-x-y)$. We plot the density function $\rho(\bm{x})$ obtained by our APRFM and reference solution in Figure 10. We set $J_{i}^{\rho}=128,J_{i}^{g}=96$ and $B=3$, with the activation function defined as $\sin(\pi x)$. The number of interior collocation points is $(N_{x_{1}},N_{x_{2}},N_{v})=(32,32,64)$, while the number of boundary collocation points is $(32,64)$. Furthermore, $16$ points are employed in the Gauss-Legendre quadrature. The domain is partitioned as $(M_{x_{1}},M_{x_{2}},M_{v})=(2,2,1)$, resulting in $M^{\rho}=4,M^{g}=4$. The relative $\ell^{2}$ error of density function $\rho(\bm{x})$ obtained by our APRFM is $4.92\times 10^{-3}(\varepsilon=1)$ and $4.09\times 10^{-3}(\varepsilon=10^{-3})$.
(a) Reference solution v.s. APRFM solution ($\varepsilon=1$).
(b) Reference solution v.s. APRFM solution ($\varepsilon=10^{-3}$).
Figure 10: Plots of $\rho(x_{1},x_{2})$ in $(x_{1},x_{2})\in[-1,1]\times[-1,1]$ by APRFM and finite difference method.
Example 5
Next, we consider a square domain with $\bm{x}=(x_{1},x_{2})\in[-1,1]^{2},\bm{v}=(\cos(\alpha),\sin(\alpha))$.
$\begin{cases}\displaystyle\varepsilon\bm{v}\cdot\nabla_{\bm{x}}f(\bm{x},\bm{v})=\frac{1}{2\pi}\int_{|\bm{v}|=1}f(\bm{x},\bm{v}^{\prime})\mathop{}\!\mathrm{d}{\bm{v}^{\prime}}-f+\varepsilon^{2}G(\bm{x},\bm{v}),\\ f(-1,x_{2},\alpha)=0,\;\alpha\in[0,\pi/2]\cup[3\pi/2,2\pi],\\ f(1,x_{2},\alpha)=0,\;\alpha\in[\pi/2,3\pi/2],\\ f(x_{1},-1,\alpha)=0,\;\alpha\in[0,\pi],\\ f(x_{1},1,\alpha)=0,\;\alpha\in[\pi,2\pi].\end{cases}$ (73)
Here, the source function $G(\bm{x},\bm{v})=1/2$. We plot the density function $\rho(\bm{x})$ obtained by our APRFM and reference solution in Figure 11. We set $J_{i}^{\rho}=64,J_{i}^{g}=128$ and $B=0.5$ for $\varepsilon=1$ and $B=1$ for $\varepsilon=10^{-1}$. The number of interior collocation points is $(N_{x_{1}},N_{x_{2}},N_{v})=(32,32,32)$, while the number of boundary collocation points is $(32,32)$. Furthermore, $24$ points are employed in the Gauss-Legendre quadrature. The domain is partitioned as $(M_{x_{1}},M_{x_{2}},M_{v})=(1,1,4)$, resulting in $M^{\rho}=1,M^{g}=4$. The relative $\ell^{2}$ error of our APRFM is $3.01\times 10^{-2}(\varepsilon=1)$ and $4.19\times 10^{-2}(\varepsilon=10^{-1})$.
(a) Reference solution v.s. APRFM solution ($\varepsilon=1$).
(b) Reference solution v.s. APRFM solution ($\varepsilon=10^{-1}$).
Figure 11: Plots of $\rho(x_{1},x_{2})$ in $(x_{1},x_{2})\in[-1,1]\times[-1,1]$ by APRFM and finite difference method.
Example 6
Finally, we focus on a hollow annular domain with $\bm{x}=(x_{1},x_{2})\in[-1,1]^{2}\setminus(-1/3,1/3)^{2},\bm{v}=(\cos(\alpha),\sin(\alpha))$.
$\begin{cases}\displaystyle\varepsilon\bm{v}\cdot\nabla_{\bm{x}}f(\bm{x},\bm{v})=\frac{1}{2\pi}\int_{|\bm{v}|=1}f(\bm{x},\bm{v}^{\prime})\mathop{}\!\mathrm{d}{\bm{v}^{\prime}}-f+\varepsilon^{2}G(\bm{x},\bm{v}),\\ f(-1,x_{2},\alpha)=e^{1-x_{2}},\,f(\frac{1}{3},x_{2},\alpha)=e^{-\frac{1}{3}-x_{2}},\;\alpha\in[0,\pi/2]\cup[3\pi/2,2\pi],\\ f(1,x_{2},\alpha)=e^{-1-x_{2}},\,f(-\frac{1}{3},x_{2},\alpha)=e^{\frac{1}{3}-x_{2}},\;\alpha\in[\pi/2,3\pi/2],\\ f(x_{1},-1,\alpha)=e^{-x_{1}+1},\,f(x_{1},\frac{1}{3},\alpha)=e^{-x_{1}-\frac{1}{3}},\;\alpha\in[0,\pi],\\ f(x_{1},1,\alpha)=e^{-x_{1}-1},\,f(x_{1},-\frac{1}{3})=e^{-x_{1}+\frac{1}{3}},\;\alpha\in[\pi,2\pi].\end{cases}$ (74)
Here, the source function $G(\bm{x},\bm{v})=(-\cos(\alpha)-\sin(\alpha))\exp(-x-y)/\varepsilon$. As before, the exact solution is given by $f_{\text{ex}}=\exp(-x-y)$. We plot the density function $\rho(\bm{x})$ obtained by our APRFM and reference solution in Figure 12. We set $J_{i}^{\rho}=128,J_{i}^{g}=96$ and $B=3$ for $\varepsilon=1$ and $B=2$ for $\varepsilon=5\times 10^{-3}$, with the activation function defined as $\sin(\pi x)$. The number of interior collocation points is $(N_{x_{1}},N_{x_{2}},N_{v})=(32,32,64)$, while the number of boundary collocation points is $(32,64)$. Furthermore, $24$ points are employed in the Gauss-Legendre quadrature. The domain is partitioned as $(M_{x_{1}},M_{x_{2}},M_{v})=(1,1,4)$, resulting in $M^{\rho}=1,M^{g}=4$. The relative $\ell^{2}$ error of our APRFM is $3.18\times 10^{-4}(\varepsilon=1)$ and $8.74\times 10^{-5}(\varepsilon=5\times 10^{-3})$.
(a) Reference solution v.s. APRFM solution ($\varepsilon=1$).
(b) Reference solution v.s. APRFM solution ($\varepsilon=5\times 10^{-3}$).
Figure 12: Plots of $\rho(x,y)$ in $(x_{1},x_{2})\in[-1,1]\times[-1,1]$ by APRFM and finite difference method.
5 Conclusions
In this paper, we proposed a novel efficient Random Feature Method based on micro-macro decomposition for efficiently solving multiscale radiative transfer equations. Motivated by a simple example, we observed that the vanilla RFM is unsuitable for the kinetic regime, while the proposed APRFM effectively addresses the RTE in this context. Extensive numerical experiments have been conducted to verify the effectiveness of the APRFM for both 1D and 2D RTEs. The numerical results of the APRFM demonstrate that it can achieve high accuracy with fewer degrees of freedom and collocation points compared with the vanilla RFM. The APRFM is robust with respect to the scale parameter $\varepsilon$ and can be applied to mixed multi-scale RTEs. The APRFM is a promising method for solving multiscale radiative transfer equations, and it has the potential to be extended to time-dependent and nonlinear kinetic equations, as well as complex geometries. The least squares problem in the APRFM, however, is more challenging to solve compared to traditional numerical methods. This involves factors such as the random initialization of network parameters, the selection of activation functions, and the choice of the PoU function, efficient sampling strategies, and domain decomposition techniques. Besides, the analysis of the convergence and stability of the APRFM is still an open question. In future work, we plan to enhance the efficiency and accuracy of the APRFM in solving kinetic equations and to extend the method to time-dependent and nonlinear kinetic equations. We also aim to investigate the convergence and stability of the APRFM and to explore the application of the APRFM to other multiscale problems.
Acknowledgments
Jingrun Chen is partially supported by the NSFC Major Research Plan - Interpretable and General Purpose Next-generation Artificial Intelligence (Nos. 92270001 and 92370205), NSFC grant 12425113, and Key Laboratory of the Ministry of Education for Mathematical Foundations and Applications of Digital Technology, University of Science and Technology of China. Zheng Ma is supported by NSFC No. 12201401, No. 92270120 and Beijing Institue of Applied Physics and Computational Mathematics funding HX02023-60.